OpenStudy (anonymous):

find dy/dx of x^2y^2+xsin(y)=1 by implicit differentiation

OpenStudy (anonymous):

\[(x^22yy'+2xy^2)+(xcosy(y')+siny(1))=0 => x^22yy'+2xy^2+xcosyy'+siny=0 => y'(x^22y+xcosy)= -siny-2xy^2=> y' = -(siny+2xy^2)/(2x^2y+xcosy)\]

OpenStudy (anonymous):

Darn

OpenStudy (anonymous):

y' = -(2xy^2+siny)/(2x^2y+xcosy) = -(2xy^2+siny)/x(2xy+cosy)

OpenStudy (anonymous):

y' = (-2xy^2 - siny) / (2x^2y + xcosy)